# Two Sides of the Same Optimization Coin: Model Degradation and Representation Collapse in Graph Foundation Models

### v2

@@ 1 Introduction
What gVQ-MAEs Limitations and How to Solve Them? Due to domain generalization conflicts, two underexplored yet interrelated pitfalls emerge in gVQ-MAEs during GFM pre-training. These pitfalls are just like two sides of the same optimization coin: Side 1 Model Degradation: The encoder and codebook often over-suppress domain-specific representations, especially for semantically conflicting inputs. Side 2 Representation Collapse: Progressive shrinkage of the latent space constrains hidden embeddings and codebook vectors to a narrow subspace [37]. This leads the decoder to over-utilize the limited embedding subset, generating low-quality reconstructed supervision and highlighting the optimization dilemma during pre-training (coin). In Sec. 3, we attribute these issues to the Information Bottlenecks and Regularization Deficits. To address them, we propose Mixture-of-Tinkers (MoT), which consists of: ❶ Information Tinker for Two Pitfalls, which utilizes an edge-wise semantic fusion strategy to enhance the encoder and employs a mixture-of-codebooks (MoC) with a tailored gated routing network. They jointly improve the information capacity of gVQ-MAE to maintain domain discriminability and representation diversity; ❷ Regularization Tinker for Optimization Coin, which utilizes the contrastive alignment and load-balancing constraint to improve Information Tinker further. They collaborate and serve as auxiliary gradient supervision.
@@ 2.1 Notations and Problem Formulation
Consider a text-attributed graph (TAG) $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\mathcal{T}\right)$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges. $\mathcal{T}$ is the textual description for nodes and edges, and the adjacency matrix is $\mathbf{A}$. To achieve GFM, $\mathcal{T}$ is encoded into feature vectors $\mathcal{X}$ using a pre-trained text encoder, where $d$ is the feature dimension. Now, we have $\mathcal{G}^{\prime}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$, which preserves both the topology and textual attributes (i.e., graph semantics). Based on this, GFM aims to learn an optimal embedding function parameterized by $f_{\theta}$ over the multi-domain $\{\mathcal{G}_{i}^{\prime}\}$ using self-supervised paradigms (e.g., mask reconstruction [9; 6] and contrastive alignment [16; 29]). Then, for cross-task requirements (e.g., node-, link-, graph-level), $f_{\theta}$ is fine-tuned with task-specific heads for better performance. In this work, we focus on multi-domain GFM pre-training and provide the detailed implementations of fine-tuning in Appendix E.3.
@@ 2.2 Graph Foundation Models
Recent advances in GFMs can be categorized into two frameworks: Language-oriented methods, which convert graphs into flattened textual representations for token encoders, and Graph-oriented methods, which preserve both textual semantics and structural integrity via dedicated architectures.
@@ 2.3 Graph Vector Quantized-Masked AutoEncoder
Most recent graph-oriented GFMs adopt gVQ-MAEs as the trainable module [32]. This architecture enables the joint encoding of topology and textual attributes into a discrete embedding space with clear semantic boundaries, making it particularly suitable for multi-domain GFM pre-training [38]. ➀ $\mathcal{G}^{\prime}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$ $\to$ Encoder $\to$ Hidden Embeddings: To ensure generality in arbitrary inputs, we propose using an encoder instantiated as any reasonable GNN capable of incorporating both node and edge features to generate informative $z\in\mathbb{R}^{d}$. ➁ Hidden Embeddings $z$ $\to$ Codebook $\to$ Quan. Emb.: To establish clear semantic boundaries, codebook $\mathcal{C}$ transforms continuous embeddings $z$ into discrete codebook vectors $e$ (equal to Quan. Emb. $z_{q}$) via similarity retrieval-based vector quantization:
@@
Quan. Emb. $z_{q}$ $\to$ Decoder $\to$ $\mathcal{G}_{r}^{\prime}=\left(\mathcal{V},\mathcal{E}_{r},\mathcal{X}_{r}\right)$: To enable self-supervised training, gVQ-MAEs follow an autoencoder framework, where gradients are computed by the discrepancy between the reconstructed supervision $\mathcal{G}_{r}^{\prime}$ and the original input $\mathcal{G}^{\prime}$, thereby updating the encoder and codebook. Notably, the trainable components of the encoder and the codebook are the weighted matrix and the discrete embeddings $\{e_{1},\dots,e_{K}\}$, which together constitute the trainable GFM embedding function parameterized by $f_{\theta}$. Meanwhile, to construct end-to-end gradient flow, the straight-through estimator (STE) [2; 35] is used to approximate gradients by passing the non-differentiable quantization step.
In gVQ-MAEs, the Side 1 Model Degradation comprises a trainable encoder and a learnable codebook. The decoder is not typically involved in fine-tuning, so we do not give it extra thought. The Side 2 Representation Collapse comprises the hidden embeddings produced by the encoder and the codebook vectors in ➁ Hidden Embeddings $z$ $\to$ Codebook $\to$ Quan. Emb.. Regarding the reconstructed supervision, as it primarily reflects the adverse effects induced by the two pitfalls (sides), we only consider it as a derivative. Besides, the optimization coin captures the overall convergence dynamics of GFM pre-training, highlighting the role of gradient-based supervision. The mask mechanisms and decoder in MoT follow conventional designs, with details provided in Sec. 4.3.
@@ 3 Empirical Investigation
As previously mentioned, model degradation and representation collapse are two sides of the same optimization coin (the convergence dilemma in multi-domain GFM pre-training). To further illustrate it, we first investigate two pitfalls (sides) via the embedding landscape (Fig. 2(a)-(c)). Then, we present the convergence curves to illustrate the direct effects of the two pitfalls in optimization (coin) (Fig. 2(d)). Please refer to Appendix B for more details about the experimental setup and analysis.
@@ 3.2 Same Coin: Optimization Dilemma in Multi-domain GFM Pre-training
Questions $\to$ Observations $\to$ Conclusions. Questions: ➀ Can gVQ-MAEs stand out? ➁ Can Information Tinker improve gVQ-MAEs? Observations: Based on the curves in Fig. 2(d), we have ➀ Compared to other GFM architectures, gVQ-MAE demonstrates its superiority; ➁ Compared to other gVQ-MAEs, Information Tinker (Naive MoT) achieves the best performance but unsatisfactory convergence. Conclusions: Although Information Tinker is effective, it remains improved.
Key Insights $\to$ Solutions: ➀ Key Insight (Sec. 5.2 ➂): The conventional gVQ-MAE commitment loss fails to effectively optimize MoC, as it merely minimizes the pairwise distances between hidden embeddings and their assigned codebook vectors, while neglecting the semantic conflicts among the codebooks. Solution: Embedding-Vector Contrastive Alignment. We pull hidden embeddings and their assigned codebook vectors closer, while incorporating the repulsion to alleviate overcrowding among MoC. This design effectively achieves Adversarial Regularization in encoder and MoC. ➁ Key Insight (Sec. 5.2 ➃): The conventional MoE load loss fails to constrain MoC, as it only enforces average expert activation without accounting for the inter-codebook preferences. In other words, rather than enforcing average codebook (expert) selection, our focus shifts toward the identification of the domain-optimal codebook. Solution: MoC Load-balancing Constraint. We dynamically redistribute MoC toward the domain-optimal load, while preventing individual codebooks from becoming high-density hubs. These two objectives collaborate and achieve Domain-aware Regularization in MoC.
@@ 4 Mixture-of-Tinkers
Figure 3: The overview of vanillin gVQ-MAE and its enhancement by our proposed MoT.
@@ 4.1 Information Tinker
Edge-wise Semantic Fusion. Existing GNNs for TAGs often naively integrate edge features by directly adding them to neighborhood aggregation outputs [25; 12]. However, when edges share identical features (i.e., encoded by a language model from raw text with the same semantics), this static approach causes information redundancy where nodes repeatedly aggregate duplicate edge features across layers. To resolve this, we propose Edge-wise Semantic Fusion, where each edge $\mathbf{e}_{uv}$ evolves by assimilating knowledge from its connected nodes $\mathbf{h}_{u}$ and $\mathbf{h}_{v}$. Formally, at layer $l$:
@@ 4.2 Regularization Tinker
4.3 Generative Self-supervision: Graph Reconstruction
where $z_{i}^{f}$ denotes the linearly projected node embeddings $z_{i}$ and $\mathbf{x}_{i}$ represents the original feature. Topology reconstruction employs a negative sampling strategy to preserve graph connectivity patterns:
@@ 4.4.1 Why Information Tinker Alleviates Semantic Entanglement?
Traditional gVQ-MAEs suffer from semantic entanglement due to ➀ Static Edge Integration: Naive aggregation schemes constrain edge-aware information flow, limiting $I(Z;\mathcal{E})$. ➁ Single-Codebook Quantization: A shared codebook $\mathcal{C}$ forces all domains into a single latent space, causing $I(Z;S)\leq\log K$ (bounded by codebook size). To break these limitations, we propose Information Tinker as described in Sec. 4.1, which comprises two key modules designed to enhance Information Flow and extend Information Resources. Based on this, we present the following theoretical foundations from an information-theoretic perspective to support their effectiveness (proofs are shown in Appendix D):
@@ 5 Experiments
To validate the superiority of MoT, we conduct comprehensive experiments. We aim to answer: Q1: Does MoT outperform SOTA baselines in supervised/few-shot/zero-shot scenarios while adhering to GFM scaling laws? Q2: How do Information and Regularization Tinker alleviate model degradation and collapse? Q3: Is MoT resilient to data-scale variations and hyper-parameter sensitivity? Q4: Does MoT achieve practical time-accuracy trade-offs? The implementation details and MoT variants are introduced in Appendix E and A.2. Additional experimental results and hyper-parameter settings can be found in Appendix F and G. Unless otherwise specified, MoT refers to the MoT-st-base.
@@ 5.1 Overall Performance
To answer Q1, we conduct a systematic evaluation of MoT across three fundamental graph learning tasks (node classification, edge classification, and graph classification) under two distinct learning paradigms: (1) supervised fine-tuning and (2) few-shot/zero-shot transfer. We compare MoT with three categories of baselines: traditional supervised GNN (GCN), unsupervised GNNs (GraphMAE, GIANT), and GFMs (GFT, OFA, SAMGPT, UniGraph).
Table 1 demonstrates MoT variants’ strong performance in supervised scenarios, where even the lightest variant MoT-st-tiny exceeds all baseline models. Compared to traditional GNNs, MoT benefits from large-scale pre-training. Compared to existing graph foundation models, MoT achieves superior generalization through its dual-tinker architecture: ➀ the information tinker dynamically fuses edge semantics to prevent model degradation, while ➁ the regularization tinker enforces geometric constraints via contrastive alignment to avert representation collapse. Table 2 highlights MoT’s exceptional adaptability in data-scarce scenarios, which consistently outperforms existing methods across few-shot and zero-shot learning settings. These results stems from the synergistic dual-tinker architecture, where the Mixture-of-Codebooks intrinsically captures transferable domain invariances, while the regularization tinker enforces geometric consistency in latent representations. Such design achieves enhanced generalization by universalizing structural patterns across domains.
Crucially, MoT demonstrates remarkable adherence to GFM scaling laws across all evaluation dimensions. As model scale increases from tiny to large variants, we observe consistent performance improvements while maintaining stable variance patterns. MoT-st-large achieves peak performance on 6 of 8 fine-tuning tasks and 15 of 16 few-shot settings, though notably underperforms MoT-st-base in certain cases. While MoT variants successfully follow scaling law, several critical issues remain unaddressed. For instance, insufficient pre-training data may lead to suboptimal parameter utilization and diminished performance returns, highlighting substantial room for future research.
@@ 5.2 Ablation Study
To address Q2, we conduct rigorous ablation studies isolating core components of MoT’s dual-tinker architecture, as shown in Table 3. We evaluate four critical variations by disabling Information Tinker (w/o. Fusion and w/o. MoC) and Regularization Tinker (w/o. $\mathcal{L}_{con}$ and w/o. $\mathcal{L}_{load}$). ➀ Edge Semantic Fusion Ablation. To evaluate the contribution of edge-wise semantic fusion, we replace it with a naive message-passing. This ablation causes significant performance degradation across all domains and tasks, which confirms that our dynamic edge feature integration is essential for addressing model degradation. ➁ Mixture-of-Codebooks Ablation. We replace the MoC with a single shared codebook architecture, which causes performance collapse, confirming its critical role in preventing representation collapse. The disentangled latent spaces by MoC preserve domain-specific semantics and capture transferable domain invariances, which are essential for cross-domain generalization. ➂ Contrastive Loss Ablation. We substitute the $\mathcal{L}_{con}$ with standard VQ loss and commitment loss [22] for codebook updates, which results in catastrophic performance degradation. By enforcing geometric separability among cross-domain representations, $\mathcal{L}_{con}$ prevents representation collapse in codebook and embedding spaces. ➃ Load Balancing Ablation. We disable $\mathcal{L}_{load}$ and use the traditional load loss [19], leading to severe routing imbalance. This proves the constraint’s necessity for balanced resource allocation across domain-specific codebooks, mitigating expert specialization bias.
@@ 5.3 Robustness Analysis
Hyper-parameter Sensitivity. We systematically analyze the impact of three key hyper-parameters: the scale of the codebook, the Top-$k$ value in the MoC, and the temperature coefficient $\tau$ in contrastive learning. As demonstrated in Fig. 5, MoT maintains robust performance across all configurations. Regarding the codebook structure, performance improves as both the codebook size and the number of codebooks increase. However, larger codebooks also introduce higher computational and memory costs. The routing mechanism and contrastive learning also play critical roles in model behavior. Excessively large or small Top-$k$ and $\tau$ can lead to performance degradation. We suggest that in the experiment, the Top-$k$ be set to 2 or 3, the $\tau$ be set to around 0.5, and the size of the codebook can be set to 16 codebooks of size 256. In this way, an optimal balance between performance and efficiency can be achieved. The recommended selection scheme of $\lambda_{1}$-$\lambda_{4}$ in Eq. 9 is described in Appendix G.1.

### v3

@@ 1 Introduction
What are gVQ-MAEs’ Limitations and How to Solve Them? Due to domain generalization conflicts, two underexplored yet interrelated pitfalls emerge in gVQ-MAEs during GFM pre-training. These pitfalls are just like two sides of the same optimization coin: Side 1 Model Degradation: The encoder and codebook often over-suppress domain-specific representations, especially for semantically conflicting inputs. Side 2 Representation Collapse: Progressive shrinkage of the latent space constrains hidden embeddings and codebook vectors to a narrow subspace [37]. This leads the decoder to over-utilize the limited embedding subset, generating low-quality reconstructed supervision and highlighting the optimization dilemma during pre-training (coin). In Sec. 3, we attribute these issues to the Information Bottlenecks and Regularization Deficits. To address them, we propose Mixture-of-Tinkers (MoT), which consists of: ❶ Information Tinker for Two Pitfalls, which utilizes an edge-wise semantic fusion strategy to enhance the encoder and employs a mixture-of-codebooks (MoC) with a tailored gated routing network. They jointly improve the information capacity of gVQ-MAE to maintain domain discriminability and representation diversity; ❷ Regularization Tinker for Optimization Coin, which utilizes the contrastive alignment and load-balancing constraint to improve Information Tinker further. They collaborate and serve as auxiliary gradient supervision.
@@ 2.1 Notations and Problem Formulation
Consider a text-attributed graph (TAG) $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\mathcal{T}\right)$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges. $\mathcal{T}$ is the textual description for nodes and edges, and the adjacency matrix is $\mathbf{A}$. To achieve GFM, $\mathcal{T}$ is encoded into feature vectors $\mathcal{X}$ using a pre-trained text encoder. Now, we have $\mathcal{G}^{\prime}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$, which preserves both the topology and textual attributes. Based on this, GFM aims to learn an optimal embedding function $f_{\theta}$ over the multi-domain $\{\mathcal{G}_{i}^{\prime}\}$ using self-supervised paradigms [9; 6; 16; 29]. For cross-task requirements, $f_{\theta}$ is fine-tuned with task-specific heads for better performance.
@@ 2.3 Graph Vector Quantized-Masked AutoEncoder
Most recent graph-oriented GFMs adopt gVQ-MAEs as the trainable module [32]. This architecture enables the joint encoding of topology and textual attributes into a discrete embedding space [38]. ➀ $\mathcal{G}^{\prime}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)\to$ Encoder $\to$ Hid. Emb.: To ensure generality, we use an encoder instantiated as any reasonable GNN capable of incorporating both node and edge features to generate $z$. ➁ Hid. Emb. $z$ $\to$ Codebook $\to$ Quan. Emb.: To establish clear semantic boundaries, codebook $\mathcal{C}$ transforms continuous $z$ into discrete codebook vectors $e$ via similarity retrieval-based vector quantization:
@@
Quan. Emb. $z_{q}$ $\to$ Decoder $\to$ $\mathcal{G}_{r}^{\prime}=\left(\mathcal{V},\mathcal{E}_{r},\mathcal{X}_{r}\right)$: To enable self-supervised training, gVQ-MAEs follow an autoencoder framework, where gradients are computed by the discrepancy between the reconstructed supervision $\mathcal{G}_{r}^{\prime}$ and the original input $\mathcal{G}^{\prime}$. To construct end-to-end gradient flow, the straight-through estimator (STE) [2; 35] is used to pass the non-differentiable quantization step.
In gVQ-MAEs, the Side 1 Model Degradation comprises a trainable encoder and codebook. The Side 2 Representation Collapse comprises the hidden embeddings and the codebook vectors. Besides, the optimization coin captures the overall GFM pre-training convergence, highlighting the role of gradient-based supervision. Our mask mechanisms and decoder are introduced in Appendix D.2.
@@ 3 Empirical Investigation
To further illustrate the model degradation and representation collapse, we first investigate two pitfalls (sides) via the embedding landscape (Fig. 2(a)-(c)). Then, we present the convergence curves to illustrate the direct effects of the two pitfalls in optimization (coin) (Fig. 2(d)). Please refer to Appendix B for more details about the experimental setup and analysis.
@@ 3.2 Same Coin: Optimization Dilemma in Multi-domain GFM Pre-training
Questions $\to$ Observations $\to$ Conclusions. Questions: ➀ Can gVQ-MAEs stand out? ➁ Can Information Tinker improve gVQ-MAEs? Observations: Based on the curves in Fig. 2(d), we have ➀ Compared to other GFM architectures, gVQ-MAE demonstrates its superiority; ➁ Compared to other gVQ-MAEs, Information Tinker (Naive MoT) achieves the best performance but unsatisfactory convergence. Conclusions: Although Information Tinker is effective, it remains to be improved.
Key Insights $\to$ Solutions: ➀ Key Insight (Sec. 5.2 ➂): The conventional gVQ-MAE commitment loss fails to effectively optimize MoC, as it merely minimizes the pairwise distances between hidden embeddings and assigned codebook vectors, while neglecting the semantic conflicts among the codebooks. Solution: Embedding-Vector Contrastive Alignment. We pull hidden embeddings and assigned codebook vectors closer, while incorporating the repulsion to alleviate overcrowding among MoC, achieving Adversarial Regularization in encoder and MoC. ➁ Key Insight (Sec. 5.2 ➃): The conventional MoE load loss fails to constrain MoC, as it only enforces average expert activation without accounting for the inter-codebook preferences. Solution: MoC Load-balancing Constraint. We dynamically redistribute MoC toward the domain-optimal load, while preventing individual codebooks from becoming high-density hubs, achieving Domain-aware Regularization in MoC.
@@ 4 Mixture-of-Tinkers
In this section, we present the details of the MoT, and Fig. 3 illustrates its complete workflow.
@@ 4.1 Information Tinker
Figure 3: The overview of vanilla gVQ-MAE and its enhancement by our proposed MoT.
Edge-wise Semantic Fusion. Existing GNNs for TAGs often naively integrate edge features by repeatedly aggregate them across layers [25; 12], which causes information redundancy when edges share identical features, mentioned in Sec. 3.1 and Appendix C. To resolve this, we propose Edge-wise Semantic Fusion, where each edge $\mathbf{e}_{uv}$ evolves by assimilating knowledge from its connected nodes $\mathbf{h}_{u}$ and $\mathbf{h}_{v}$. The detailed operation are described in Appendix D.1.
$\mathbf{h}_{u}^{(l+1)}=\operatorname{Agg}_{1}\left(\mathbf{h}_{u}^{(l)};\operatorname{Prop}_{1}\left\{\mathbf{h}_{v}^{(l)},\mathbf{e}_{uv}^{(l)}\right\}\right),\;\;\mathbf{e}_{uv}^{(l+1)}=\operatorname{Agg}_{2}\left(\mathbf{e}_{uv}^{(l)};\operatorname{Prop}_{2}\left\{\mathbf{h}_{u}^{(l)},\mathbf{h}_{v}^{(l)}\right\}\right).$ (2)
@@ 4.3.1 Why Information Tinker Alleviates Semantic Entanglement?
Traditional gVQ-MAEs suffer from semantic entanglement due to ➀ Static Edge Integration: Naive aggregation schemes constrain edge-aware information flow, limiting $I(Z;\mathcal{E})$. ➁ Single-Codebook Quantization: A shared codebook $\mathcal{C}$ forces all domains into a single space, causing $I(Z;S)\leq\log K$ (bounded by codebook size). To break these limitations, we propose Information Tinker and present the following theoretical foundations to support its effectiveness (proofs are shown in Appendix E):
@@ 5 Experiments
To validate the superiority of MoT, we conduct comprehensive experiments. We aim to answer: Q1: Does MoT outperform SOTA baselines in supervised/few-shot/zero-shot scenarios while adhering to GFM scaling laws? Q2: How do Information and Regularization Tinker alleviate model degradation and collapse? Q3: Is MoT resilient to data-scale variations and hyper-parameter sensitivity? Q4: Does MoT achieve practical time-accuracy trade-offs? The implementation details and MoT variants are introduced in Appendix F and A.2. Additional experimental results and hyper-parameter settings can be found in Appendix G and H. Unless otherwise specified, MoT refers to the MoT-st-base.
@@ 5.1 Overall Performance
To answer Q1, we conduct systematic evaluations across three fundamental graph learning tasks (node/edge/graph classifications) under two distinct learning paradigms: (1) supervised fine-tuning and (2) few-/zero-shot transfer. We compare MoT with three categories of baselines: supervised GNN (GCN), unsupervised GNNs (GraphMAE, GIANT), and GFMs (GFT, OFA, SAMGPT, UniGraph).
Table 1 demonstrates MoT variants’ strong performance in supervised scenarios, where even the lightest variant MoT-st-tiny exceeds all baselines. Table 2 highlights MoT’s adaptability in data-scarce scenarios, which consistently outperforms existing methods across few-/zero-shot learning settings. Compared to traditional GNNs, MoT benefits from large-scale pre-training. Compared to existing GFMs, MoT achieves superior generalization through its dual-tinker architecture: ➀ the Information Tinker dynamically fuses edge semantics to prevent model degradation, and ➁ the Regularization Tinker enforces geometric constraints via contrastive alignment to avert representation collapse. Such design achieves enhanced generalization by universalizing structural patterns across domains.
Crucially, MoT demonstrates remarkable adherence to GFM scaling laws across all evaluation dimensions. As model scale increases from tiny to large variants, we observe consistent performance improvements while maintaining stable variance patterns. MoT-st-large achieves peak performance on 6 of 8 fine-tuning tasks and 15 of 16 few-shot settings, though notably underperforms MoT-st-base in certain cases. While MoT variants successfully follow the scaling law, several critical issues remain unaddressed. For instance, insufficient pre-training data may lead to suboptimal parameter utilization and diminished performance returns, highlighting substantial room for future research.
@@ 5.2 Ablation Study
To address Q2, we conduct ablation studies isolating core components of MoT, as shown in Table 3. We evaluate four critical variations by disabling Information Tinker (w/o. Fusion and w/o. MoC) and Regularization Tinker (w/o. $\mathcal{L}_{con}$ and w/o. $\mathcal{L}_{load}$). ➀ Edge Semantic Fusion Ablation. We replace the edge-wise semantic fusion with a naive message-passing, which causes significant performance degradation across all domains and tasks, confirming that our dynamic edge feature integration is essential for addressing model degradation. ➁ MoC Ablation. We replace the MoC with a single codebook, which causes performance collapse, confirming its critical role in preventing representation collapse. MoC preserves domain semantics and captures transferable domain invariances, which are essential for cross-domain generalization. ➂ Contrastive Loss Ablation. We substitute the $\mathcal{L}_{con}$ with standard commitment loss [22] for codebook updates, which results in catastrophic performance degradation. By enforcing geometric separability among cross-domain representations, $\mathcal{L}_{con}$ prevents representation collapse in codebook and embeddings. ➃ Load Balancing Ablation. We disable $\mathcal{L}_{load}$ and use the traditional load loss [19], leading to severe routing imbalance. This proves the constraint’s necessity for balanced resource allocation in MoC, mitigating expert specialization bias.
@@ 5.3 Robustness Analysis
Hyper-parameter Sensitivity. We systematically analyze the impact of three key hyper-parameters: the scale of the codebook, the Top-$k$ value in the MoC, and the temperature coefficient $\tau$ in contrastive learning. As demonstrated in Fig. 5, MoT maintains robust performance across all configurations. Regarding the codebook structure, performance improves as both the codebook size and the number of codebooks increase. However, larger codebooks also introduce higher computational and memory costs. The routing mechanism and contrastive learning also play critical roles in model behavior. Excessively large or small Top-$k$ and $\tau$ can lead to performance degradation. We suggest that in the experiment, the Top-$k$ be set to 2 or 3, the $\tau$ be set to around 0.5, the size of the codebook be set to 16 codebooks of size 256, which achieves an optimal balance between performance and efficiency.